OpenStudy (anonymous):

How many combinations are possible with repeats for C(5, 2)

OpenStudy (anonymous):

@SolomonZelman Could you please help me?

OpenStudy (solomonzelman):

\(\LARGE\color{midnightblue}{ \rm n~C~r~=~\frac{n!}{r!(n-r)!} }\)

OpenStudy (solomonzelman):

\(\LARGE\color{purple}{ \rm 5~C~2~=~\frac{5!}{2!(5-2)!} }\)

OpenStudy (solomonzelman):

Finish the calculations from here.

OpenStudy (anonymous):

What does the "!" symbol mean?

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm 8!=1 \times 2 \times 3 \times 4 \times 5 \times 6 \times 7 \times 8 }\) \(\large\color{green}{ \rm 7!=1 \times 2 \times 3 \times 4 \times 5 \times 6 \times 7 }\) \(\large\color{red}{ \rm 6!=1 \times 2 \times 3 \times 4 \times 5 \times 6 }\) \(\large\color{blue}{ \rm 5!=1 \times 2 \times 3 \times 4 \times 5 }\) \(\large\color{darkorange }{ \rm ~~BUT,~~~ 0!=1~~~(exception) }\)

OpenStudy (solomonzelman):

See the idea behind the ` ! ` ?

OpenStudy (imer):

Its mean "factorial", For example 5! = 5*4*3*2*1

OpenStudy (anonymous):

Thank you!